### v2

# Ricci curvature, diameter and eigenvalues of amply regular graphs

Abstract
We prove a weaker version of a conjecture proposed by Qiao, Park and Koolen on diameter bounds of amply regular graphs and make new progress on Terwilliger’s conjecture on finiteness of amply regular graphs. We achieve these results by a significantly improved Lin–Lu–Yau curvature estimate and new Bakry–Émery curvature estimates. We further discuss applications of our curvature estimates to bounding eigenvalues, isoperimetric constants, and volume growth. In particular, we obtain a volume estimate of amply regular graphs which is sharp for hypercubes.
1 Introduction
It was one of the main open problems for distance regular graphs to show that there are only finitely many distance-regular graphs with fixed vertex degree $d\geq 3$ [45, 46]. This was later known as the Bannai-Ito conjecture, referring to their 1984 book [5, p.237] and has been confirmed in 2015 by Bang, Dubickas, Koolen and Moulton [4]. The study of Bannai-Ito conjecture has particularly stimulated a very fruitful approach of bounding the diameter of a distance regular graph in terms of its vertex degree, see, e.g., [10, 18, 31, 34, 35, 43, 45, 46].
During the process of investigating the diameter bounds of distance regular graphs, it was realized that some diameter bounds hold even for graphs with only a fraction of the regularity properties distance-regular graphs possess. For example, Mulder [31] obtained in 1979 a diameter bound for $(0,\lambda)$-graphs with $\lambda\geq 2$. Recall that a $(0,\lambda)$-graph is a connected simple graph such that any two distinct vertices have either 00 or $\lambda$ common neighbors. Terwilliger [46] derived in 1983 a diameter bound for $(s,c,a,k)$-graphs with integers $s,c,a+2,k\geq 2$. In case that $s=2$, a $(2,c,a,k)$-graph is a connected amply regular graph with parameter $(n,k,a,c)$, i.e., a $d$-regular graph such that any two adjacent vertices have $a$ common neighbors and any two vertices at distance $2$ have $c$ common neighbors. The parameter $n$ is the number of vertices. We allow $n=\infty$ here. Recently, Neumaier and Penjić [35] raised the question of bounding the diameter of a distance regular graph with a small initial part of its intersection array. In particular, they asked for diameter bounds for a distance regular graph in terms of the first four number $\{b_{0},b_{1};c_{1},c_{2}\}$ of its intersection array. Notice that $c_{1}=1$ and a distance regular graph is an amply regular graph with parameter $(n,b_{0},b_{0}-b_{1}-1,c_{2})$.
In this paper we study diameter bounds for connected amply regular graphs with parameters $(n,d,\alpha,\beta)$. The number of vertices $n$, vertex degree $d$ and parameters $\alpha,\beta$ are nonnegative integers, with $n$ being possibly infinity. For convenience, we exclude the complete graphs and empty graphs. In particular, we then have $n,d,\beta>0$. Below we recall briefly several known diameter bounds and related open questions for an amply regular graph $G$. Inspired by a diameter bound due to Mulder for $(0,\lambda)$-graphs, Brouwer, Cohen and Neumaier [10, Corollary 1.9.2] prove that
$\mathrm{diam}(G)\leq\max\{4,d-2\beta+4\},\,\,\text{if}\,\,1\neq\beta\geq\alpha.$ (1.1)
Terwilliger’s diameter bounds in [46, Theorem 3.4] imply particularly that
$\mathrm{diam}(G)\leq d,\,\,$ if  $\text{if}\,\,2=\beta\geq\alpha,$ (1.2)
$\mathrm{diam}(G)\leq\max\left\{5,\frac{2\beta(d-1)}{3\beta-2}-2\beta+7\right\},\,\,$ $\text{if}\,\,\beta\geq\max\{3,\alpha\}.$ (1.3)
Notice that the diameter bound (1.3) is stronger than (1.1) for large $d$. Particularly, we point out that the coefficient of $d$ in (1.3) satisfies $1>\frac{2\beta}{3\beta-2}>\frac{2}{3}$. Terwilliger [46, p.160] further proposed the following conjecture.
Conjecture 1.1 (Terwilliger).
Any amply regular graph with parameters $(n,d,\alpha,\beta)$ where $\beta\geq 2$ is always finite, i.e., $n<\infty$.
We mention that both the diameter bounds (1.2)-(1.3) and the conjecture of Terwilliger are made for general $(s,c,a,k)$-graphs. We reduce them to amply regular graphs here. Terwilliger’s diameter bounds actually confirm Conjecture 1.1 for amply regular graphs satisfying $1\neq\beta\geq\alpha$. Notice that amply regular graphs with $\beta=1$ can be infinite. Restricting to the class of finite graphs, there still exist amply regular graphs with $\beta=1$ and fixed degree, whose number of vertices can be arbitrarily large, see Appendix for an explicit construction. A related natural question is whether or not there are only finitely many amply regular graphs with $\beta\geq 2$ and fixed degree.
Qiao, Park, and Koolen [37] observed from Terwilliger’s diameter bounds that any connected amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta>\frac{d}{3}\geq 1$ and $\beta\geq\alpha$ has diameter at most $7$. Based on this phenomenon, they [37, Conjecture 1.2] proposed the following conjecture.
Conjecture 1.2 (Qiao, Park and Koolen).
Let $G$ be a connected amply regular graph with parameters $(n,d,\alpha,\beta)$ and $10<\varepsilon<$ 1 be a real number. Then there exists a constant $K(\varepsilon)$ such that, if $\beta>\varepsilon d\geq 1$ and $\beta\geq\alpha$ all hold, then the diameter of $G$ is at most $K(\varepsilon)$.
Recently, Huang, the third named author and Xia [22] show for an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$ that
$\mathrm{diam}(G)\leq\left\lfloor\frac{2d}{3}\right\rfloor,\,\,\text{if}\,\,\beta>\alpha\geq 1.$ (1.4)
In this paper, we establish the following diameter bound for amply regular graphs.
Theorem 1.3.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Then the diameter of the graph satisfies
$\mathrm{diam}(G)\leq\left\lfloor\frac{2d}{2+\max\left\{\frac{\alpha}{2},\left\lceil\frac{\beta-\alpha}{\beta-1}\alpha\right\rceil\right\}}\right\rfloor,\,\,\text{if}\,\,\beta\geq\max\{3,\alpha\}.$ (1.5)
We remark that (1.5) also holds for the case $2=\beta>\alpha$. That is, (1.5) actually holds for the case $1\neq\beta\geq\alpha$ except for $\alpha=\beta=2$. In case $\alpha=\beta=2$, both our method and the existing bound (1.1) and (1.2) can only tell $\mathrm{diam}(G)\leq d$ (see Corollary 4.5 below). We wonder whether or not the bound (1.5) still holds for amply regular graphs with $\alpha=\beta=2$.
Theorem 1.3 improves previous diameter bounds (1.1),(1.3) and (1.4) (for large $d$). Indeed, the coefficient of $d$ in our diameter bound (1.5) can be much smaller than $\frac{2}{3}$. In particular, if $\beta>\alpha^{2}-\alpha+1$, then Theorem 1.3 tells $\mathrm{diam}(G)\leq\left\lfloor\frac{2}{2+\alpha}d\right\rfloor$. We refer to Corollary 3.5 for the existence of a local perfect matching in this case.
Moreover, Theorem 1.3 confirms a weaker version of Conjecture 1.2.
Corollary 1.4.
Let $G$ be a connected amply regular graph with parameters $(n,d,\alpha,\beta)$ and $10<\varepsilon<$ 1 be a real number. If $\beta\geq\alpha>\varepsilon d\geq 1$, then the diameter of $G$ is at most $4/\varepsilon$.
We prove Theorem 1.3 by discrete Ricci curvature estimates of amply regular graphs. Recall that the diameter bound (1.4) is obtained via estimating the Lin–Lu–Yau curvature [22, Theorem 1.3]. The proof of Theorem 1.3 is built upon two main ingredients: a significant improvement of the Lin–Lu–Yau curvature estimate and a formula to calculate the Bakry–Émery curvature of amply regular graphs; The former is related to local matching conditions and is achieved by using König theorem for regular bipartite graphs with possibly multiple edges, while the latter is related to the adjacency eigenvalue of the local graph induced by neighbors of a vertex and is a generalization of previous results for strongly regular graphs in the arXiv version of [16].
Lin–Lu–Yau curvature [25, 36] and Bakry–Émery curvature [3, 26] of graphs are discrete analogues of Ricci curvature (lower bounds) of Riemannian manifolds, based on optimal transportation and Bochner identity, respectively. The study of discrete Ricci curvatue has attracted lots of attention in recent years, see, e.g., the book [33] and references therein. Various combinatorial, geometric and analytic consequences of Lin–Lu–Yau and Bakry–Émery curvature lower bounds have been established, see, e.g., [7, 13, 19, 21, 23, 30, 27, 32, 38, 39, 41, 44]. Especially, positive Lin–Lu–Yau or Bakry–Émery curvature lower bound implies the finiteness of a locally finite graph and a sharp diameter bound [25, 28]. Such results are considered as discrete analogues of Bonnet-Myers theorem from Riemannian geometry.
Another interesting consequence of our discrete Ricci curvature approach is an improvement of Terwilliger’s finiteness result for amply regular graphs, concerning his Conjecture 1.1.
Theorem 1.5.
Any amply regular graph with parameters $(n,d,\alpha,\beta)$ with $\alpha\leq 6\beta-9$ is finite, i.e., $n<\infty$. Moreover, there are only finitely many amply regular graphs with $\alpha\leq 6\beta-9$ and fixed degree.
Recall that Terwilliger [46] has shown the above result for the case $1\neq\beta\geq\alpha$. Theorem 1.5 is proved via Bakry-Émery curvature lower bound estimates not only for the case $\beta\geq\alpha$ but also for certain case with $\alpha>\beta$. For Lin–Lu–Yau curvature, we obtain a positive lower bound in case $\alpha>\beta$ and $d\leq\alpha+\beta$ in Theorem 3.7.
Our discrete Ricci curvature approach further leads to spectral gap estimates for amply regular graphs. Let $\theta_{1}\leq\cdots\leq\theta_{n-1}\leq\theta_{n}$ be the adjacency eigenvalues of a $d$-regular graph $G$ with $n$ vertices. Then we have
$-d\leq\theta_{1}\leq\cdots\leq\theta_{n-1}\leq\theta_{n}=d,$
where $\theta_{n-1}=d$ if and only if $G$ is disconnected, and $\theta_{1}=-d$ if and only if $G$ is bipartite.
Theorem 1.6.
Let $G$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$.
  • (i)
    If $1\neq\beta\geq\alpha$ and $(\alpha,\beta)\neq(2,2)$, then
    $\theta_{n-1}\leq d-2-\max\left\{\frac{\alpha}{2},\left\lceil\frac{(\beta-\alpha)\alpha}{\beta-1}\right\rceil\right\}.$
  • (ii)
    If $\alpha=\beta=2$, then
    $\theta_{n-1}\leq d-2.$
This improves the previous sharp eigenvalue estimates in [22, Corollary 1.6] significantly. Notice that $n<\infty$ in Theorem 1.6 is guaranteed by the parameter conditions due to Theorem 1.5.
Moreover, our approach also produces an estimate for the gap between $\theta_{1}$ and $-d$ in the other end.
Theorem 1.7.
Let $G$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ of diameter at least $4$. If $122\leq\alpha\leq 10\beta-$ 12, then
$\theta_{1}\geq-d+2.$
We achieve this via establishing a formula to calculate the Bakry–Émery curvature of a signed graph $(G,\sigma)$ with $G$ being an amply regular graph and $\sigma$ being the all-$-1$ signature. A signed graph is a graph $G=(V,E)$ associated with a signature $\sigma:E\to\{+1,-1\}$ [17, 48]. Bakry–Émery curvature for general connection graphs, including signed graphs as a special case, is introduced by the third named author, Münch and Peyerimhoff [29], and further studied in [11, 20].
Our eigenvalue estimates in Theorem 1.6 and Theorem 1.7 naturally yields estimates on Cheeger constant, bipartiteness constant [6, 47], and expansion constant [2, p.151] of an amply regular graph, see Corollary 5.7 and Theorem 5.10.
We further derive the following sharp volume growth estimate for amply regular graphs by combining our discrete curvature estimates with two key lemmas — [25, Lemma 4.4] and [46, Lemma 3.3], which are due to Lin–Lu–Yau and Terwilliger, respectively.
Theorem 1.8.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $1\neq\beta\geq\alpha$. Then we have for any $x\in V$ and $i\geq 1$ that
$|S_{i+1}(x)|\leq\frac{d-\max\left\{\alpha+1,i\left(1+\frac{1}{2}\left\lceil\frac{\alpha(\beta-\alpha)}{\beta-1}\right\rceil\right)\right\}}{\beta+i-1}|S_{i}(x)|,$
where $S_{i}(x)$ is the set of vertices of distance $i$ to $x$.
The volume growth estimates in Theorem 1.8 and the corresponding volume estimate are sharp for hypercube graphs (see Example 5.15).
Finally, we comment that our diameter and eigenvalue estimates in Theorems 1.3, 1.6 and 1.7 hold still under more general conditions on $\alpha,\beta$ involving the degree $d$, see Corollary 4.7 and Theorem 5.1.
The rest of our paper is organized as follows. In Section 2, we provide basics about the two discrete Ricci curvature notions and about signed and amply regular graphs. We establish the Lin–Lu–Yau curvature and Bakry–Émery curvature estimates of amply regular graphs with applications to diameter bounds in Sections 3 and 4, respectively. We discuss further applications of our curvature estimates in bounding eigenvalues, isoperimetric constants and volume growth of amply regular graphs in Section 5.
2 Preliminaries
In this section, we collect preliminaries about Lin–Lu–Yau curvature, Bakry–Émery curvature, signed graphs and amply regular graphs. Let $G=(V,E)$ be a graph with vertex set $V$ and edge set $E$. For any two vertices $x,y\in V$, we write $xy\in E$ if there exists an edge $\{x,y\}\in E$ between them. The vertex degree $d_{x}$ of a vertex $x\in V$ is defined by $d_{x}=\sum_{y\in V:\,xy\in E}1$. A graph $G$ is called locally finite if $d_{x}<\infty$ for any $x\in V$.
2.1 Lin–Lu–Yau Curvature
Definition 2.1 (Wasserstein distance).
Let $G=(V,E)$ be a locally finite graph, $\mu_{1}$ and $\mu_{2}$ be two probability measures on $V$. The Wasserstein distance $W_{1}(\mu_{1},\mu_{2})$ between $\mu_{1}$ and $\mu_{2}$ is defined as
$W_{1}(\mu_{1},\mu_{2})=\inf_{\pi}\sum_{y\in V}\sum_{x\in V}d(x,y)\pi(x,y),$
where $d(x,y)$ is the combinatorial distance between $x$ and $y$ in $G$, and the infimum is taken over all maps $\pi:V\times V\to[0,1]$ satisfying
$\mu_{1}(x)=\sum_{y\in V}\pi(x,y),\,\,\mu_{2}(y)=\sum_{x\in V}\pi(x,y).$
Such a map is called a transport plan.
For any $p\in[0,1]$, consider the particular probability measure $\mu_{x}^{p}$ around a vertex $x\in V$ defined as follows:
$\mu_{x}^{p}(y)=\left\{\begin{array}[]{ll}p,&\hbox{if $y=x$;}\\ \frac{1-p}{d_{x}},&\hbox{if $xy\in E$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
The following definition of discrete curvature is introduced by Ollivier [36] and Lin, Lu and Yau [25], as a discrete analgoue of the Ricci curvature of a Riemannian manifold.
Definition 2.2 ($p$-Ollivier curvature and Lin–Lu–Yau curvature [25, 36]).
Let $G=(V,E)$ be a locally finite graph. For any vertices $x,y\in V$ and $p\in[0,1]$, the $p$-Ollivier curvature $\kappa_{p}(x,y)$ is defined as
$\kappa_{p}(x,y):=1-\frac{W_{1}(\mu_{x}^{p},\mu_{y}^{p})}{d(x,y)}.$
We then define the Lin–Lu–Yau curvature $\kappa_{LLY}(x,y)$ as
$\kappa_{LLY}(x,y):=\lim_{p\to 1}\frac{\kappa_{p}(x,y)}{1-p}.$
Notice that $\kappa_{1}(x,y)$ is always 00. Hence, the Lin–Lu–Yau curvature $\kappa_{LLY}(x,y)$ equals to the negative of left derivative of the function $p\mapsto\kappa_{p}(x,y)$ at $p=1$.
For regular graphs, Bourne et. al. [9] found the following limit-free reformulation of the Lin–Lu–Yau curvature, by showing that the curvature function $p\mapsto\kappa_{p}(x,y)$ at any edge $xy\in E$ is piecewise linear with at most $2$ linear parts.
Theorem 2.3 ([9]).
Let $G=(V,E)$ be a $d$-regular locally finite graph. For any edge $xy\in E$, we have
$\kappa_{LLY}(x,y)=2\kappa_{\frac{1}{2}}(x,y)=\frac{d+1}{d}\kappa_{\frac{1}{d+1}}(x,y).$
We have the following Lin–Lu–Yau curvature upper bound and an interesting connection to the existence of a local perfect matching, see, e.g., [13, Propsition 2.7]. The upper bound estimate has been established in [25, Lemma 4.4], and its relations with local matchings have been discussed in [8, 32, 44].
Theorem 2.4 ([13, Propsition 2.7]).
Let $G=(V,E)$ be a $d$-regular graph. For any edge $xy\in E$, we have
$\kappa_{LLY}(x,y)\leq\frac{2+|\Delta_{xy}|}{d},$
where $\Delta_{xy}:=S_{1}(x)\cap S_{1}(y)$ and $S_{1}(v):=\{w\in V|\,vw\in E\}$. Moreover, the following two propositions are equivalent:
  • (a)
    $\kappa_{LLY}(x,y)=\frac{2+|\Delta_{xy}|}{d}$;
  • (b)
    there is a perfect matching between $N_{x}$ and $N_{y}$, where $N_{x}:=S_{1}(x)\backslash(\Delta_{xy}\cup\{y\})$ and $N_{y}:=S_{1}(y)\backslash(\Delta_{xy}\cup\{x\})$.
Next, we recall Hall’s marriage theorem on the existence of a perfect matching in a bipartite graph, and its natural consequence which is result of König.
Theorem 2.5 (Hall’s marriage theorem).
Let $H=(V,E)$ be a bipartite graph with the bipartition $V=V_{1}\sqcup V_{2}$. Then $H$ has a perfect matching if and only if
$|V_{1}|=|V_{2}|\,\,\text{and}\,\,|\Gamma_{V_{2}}(S)|\geq|S|\,\,\text{for all}\,\,S\subseteq V_{1},$
where $\Gamma_{V_{2}}(S):=\{v\in V_{2}|\,\,\text{there exists }\,w\in S\,\text{such that}\,\,vw\in E\}$.
Theorem 2.6 (König’s theorem).
A bipartite graph $G$ can be decomposed into $d$ edge-disjoint perfect matchings if and only if $G$ is $d$-regular.
Notice that in Theorem 2.5 and hence Theorem 2.6, the bipartite graph is allowed to have multiple edges.
Let us conclude this subsection by recalling the following diameter estimate, which is an analogue of the Bonnet–Myers theorem in Riemannian geometry.
Theorem 2.7 (Discrete Bonnet–Myers theorem via Lin–Lu–Yau curvature [25, 36]).
Let $G=(V,E)$ be a locally finite connected graph. If $\kappa_{LLY}(x,y)\geq k>0$ holds true for any edge $xy\in E$, then the graph is finite and the diameter satisfies
$\mathrm{diam}(G)\leq\frac{2}{k}.$
2.2 Bakry–Émery curvature of signed graphs
A signed graph $(G,\sigma)$ is a graph $G=(V,E)$ associated with a signature $\sigma:E\rightarrow\{+1,-1\}$. The signature of a cycle in $G$ is defined to be the product of all signatures on edges of this cycle. A signed graph $(G,\sigma)$ is called balanced if the signature of each cycle of $G$ is $+1$, and anti-balanced if the signature of each even cycle is $+1$ and of each odd cycle is $-1$ [17].
An important characterization of balanced and anti-balanced signatures is given via the following concept of switching equivalence.
Definition 2.8 (Switching equivalence).
Let $G=(V,E)$ be a graph. Two signatures $\sigma,\sigma^{\prime}$ of $G$ are said to be switching equivalent if there exists a function $\tau:V\rightarrow\{+1,-1\}$, called a switching function, such that for each edge $xy\in E$,
$\sigma^{\prime}(xy)=\tau(x)\sigma(xy)\tau(y).$
Indeed, a signed graph $(G,\sigma)$ is balanced if and only if the signature $\sigma$ is switching equivalent to the all-$+1$ signature and is anti-balanced if and only if the signature is switching equivalent to the all-$-1$ signature [48].
For an edge $xy\in E$, We will write $\sigma_{xy}=\sigma(xy)$ for short.
Next, we recall the connection Laplacian and the corresponding Bakry–Émery curvature on a signed graph [3, 29]. The connection Laplacian $\Delta^{\sigma}$ is defined as below. For any function $f:V\rightarrow\mathbb{R}$ and any vertex $x\in V$,
$\Delta^{\sigma}f(x):=\sum_{y:\,xy\in E}(\sigma_{xy}f(y)-f(x)).$
In matrix form, we have $\Delta^{\sigma}=A^{\sigma}-D$, where $A^{\sigma}$ is the signed adjacency matrix and $D$ is the diagonal degree matrix. In case that $\sigma\equiv+1$ is the all-$+1$ signature, we write $\Delta$ instead of $\Delta^{\sigma}$ for short.
Following [29], we define the $\Gamma^{\sigma}$-operator and $\Gamma^{\sigma}_{2}$-operator as follows. For any functions $f,g:V\rightarrow\mathbb{R}$ and any vertex $x\in V$, we have
$2\Gamma^{\sigma}(f,g)(x):=$ $\Delta(fg)(x)-(\Delta^{\sigma}f)g(x)-f(\Delta^{\sigma}g)(x).$
$2\Gamma_{2}^{\sigma}(f,g)(x):=$ $\Delta\Gamma^{\sigma}(f,g)(x)-\Gamma^{\sigma}(\Delta^{\sigma}f,g)(x)-\Gamma^{\sigma}(f,\Delta^{\sigma}g)(x).$
Notice in the above definitions that both the connection Laplacians $\Delta^{\sigma}$ for the signature $\sigma$ and $\Delta$ for the all-$+1$ signature are involved. For simplicity, we denote
$\Gamma^{\sigma}(f)(x):=\Gamma^{\sigma}(f,f)(x),\,\,\Gamma^{\sigma}_{2}(f)(x):=\Gamma^{\sigma}_{2}(f,f)(x).$
Definition 2.9 (Bakry–Émery curvature [3, 29]).
Let $(G,\sigma)$ be a signed graph. We say a vertex $x$ satisfies the Bakry–Émery curvature dimension inequality $CD^{\sigma}(K,N)$ for $K\in\mathbb{R}$ and $N\in(0,\infty]$, if it holds for any function $f:V\rightarrow{\mathbb{R}}$ that
$\Gamma_{2}^{\sigma}(f)(x)\geq\frac{1}{N}(\Delta^{\sigma}f)^{2}(x)+K\Gamma^{\sigma}(f)(x).$ (2.1)
For $N\in(0,\infty]$, we define the $N$-Bakry–Émery curvature at a vertex $x$ of $(G,\sigma)$ as
$K_{BE}(G,\sigma,x,N):=\sup\{K:CD^{\sigma}(K,N)\text{ holds at }x\}.$
We are particularly interested in the case $N=\infty$, for which we use the convention that $\frac{1}{N}=0$.
Proposition 2.10 (Switching invariance [29]).
Let $(G,\sigma)$ be a signed graph with $G=(V,E)$. Let $\tau:V\to\{+1,-1\}$ be a switching function and $\sigma^{\tau}$ be the signature such that $\sigma^{\tau}_{xy}:=\tau(x)\sigma_{xy}\tau(y)$ for any $xy\in E$. Then $(G,\sigma)$ satisfy $CD^{\sigma}(K,N)$ at $x$ if and only if $(G,\sigma^{\tau})$ satisfy $CD^{\sigma^{\tau}}(K,N)$ at $x$. That is, we have
$K_{BE}(G,\sigma,x,N)=K_{BE}(G,\sigma^{\tau},x,N).$
In this paper, we are concerned with the $\infty$-Bakry–Émery curvature of a graph associated with either the all-$+1$ or the all-$-1$ signature. We will write the corresponding Bakry–Émery curvature as
$K_{BE}(+,x):=K(G,+,x,\infty),\,\,\text{and}\,\,K_{BE}(-,x):=K_{BE}(G,-,x,\infty),$
for short.
The following diameter estimate via Bakry–Émery curvature is another analogue of the Bonnet-Myers theorem in Riemannian geometry.
Theorem 2.11 (Discrete Bonnet–Myers theorem via Bakry–Émery curvature [28]).
Let $G=(V,E)$ be a locally finite connected graph. If $K_{BE}(+,x)\geq K>0$ holds true for any vertex $x\in V$, then the graph is finite and the diameter satisfies
$\mathrm{diam}(G)\leq\frac{2d_{\max}}{K},$
where $d_{\max}$ stands for the maximal vertex degree.
Remark 2.12.
We point out that similar diameter estimate as above using Bakry–Émery curvature of a general signature $\sigma$ dose NOT hold any more. In [20, Example 8.1], an anti-balanced regular graph $G=(V,E)$ with positive curvature $K_{BE}(-,x)=\frac{7-\sqrt{17}}{4}>0$ at each vertex $x\in V$ and diameter $\text{diam}(G)=\infty$ is constructed. On the other hand, Bakry–Émery curvature on general signed graphs is quite useful for studying eigenvalue estimates and heat semigroup properties [29].
The two discrete Ricci curvature notions, Bakry–Émery curvature and Lin–Lu–Yau curvature, are naturally related. However, they can also be very different: the former is defined on vertices while the later is defined on edges. There are special cases in which the two curvature values have opposite signs. We refer the readers to [14] for more comparison results between these two curvature notions.
2.3 Amply regular graphs
We recall the definition of an amply regular graph form [10, Section 1.1]. We exclude the complete and empty graphs for convenience and allow infinite graphs.
Definition 2.13 (Amply regular graph [10]).
[…]
In this paper, we restrict ourselves to connected amply regular graphs. Since we have excluded complete graphs in the definition of amply regular graphs, we naturally have $d\geq\alpha+2$. The following degree estimates will also be useful for our later purpose.
Theorem 2.14 ([10, Theorem 1.2.3]).
Let $G$ be a connected amply regular graph with parameters $(n,d,\alpha,\beta)$, then we have
$d\geq 2\alpha-\beta+3.$ (2.2)
Equality holds in (2.2) if and only if $G$ is the icosahedron or the line graph of a regular graph of girth at least $5$.
Corollary 2.15 ([10, Corollary 1.2.4]).
Let $G$ be a connected amply regular graph with parameters $(n,d,\alpha,\beta)$. If $G$ has diameter at least three and $d<3(\alpha-\beta+2)$, then equality holds in (2.2).
Theorem 2.16 ([10, Theorem 1.9.3]).
Let $G$ be a connected amply regular graph of diameter at least $4$ with parameters $(n,d,\alpha,\beta)$, then $d\geq 2\beta$.
3 Lin–Lu–Yau curvature of amply regular graphs
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. For any vertex $v\in V$, we set $S_{1}(v):=\{w\in V|\,vw\in E\}$. In this section, we fix an edge $xy\in E$ and use the notations:
$\Delta_{xy}:=S_{1}(x)\cap S_{1}(y),\ N_{x}:=S_{1}(x)\backslash(\Delta_{xy}\cup\{y\})\ {\rm and}\ N_{y}:=S_{1}(y)\backslash(\Delta_{xy}\cup\{x\}).$
It follows by definition that $|\Delta_{xy}|=\alpha\ {\rm and}\ |N_{x}|=|N_{y}|=d-\alpha-1$.
3.1 Case of $1\neq\beta\geq\alpha$
We prove the following result in this subsection.
Theorem 3.1.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $1\neq\beta\geq\alpha$. Then the Lin–Lu–Yau curvature $\kappa_{LLY}(x,y)$ of any edge $xy\in E$ satisfies
$\kappa_{LLY}(x,y)\geq\frac{2+\left\lceil\frac{\alpha(\beta-\alpha)}{\beta-1}\right\rceil}{d}.$ (3.1)
As a preparation for the proof of Theorem 3.1, we construct from the local structure of the edge $xy\in E$ a bipartite multigraph $H_{G}=(V_{H},E_{H})$ as follows. Let us denote the $\alpha$ vertices in $\Delta_{xy}$ by $z_{1},\cdots,z_{\alpha}$. The vertex set of $H$ is given by
$V_{H}=N_{x}\cup N_{y}\cup\Delta_{xy}\cup\Delta^{\prime}_{xy}.$
Here $\Delta^{\prime}_{xy}:=\{z^{\prime}_{1},\cdots,z^{\prime}_{\alpha}\}$ is a new added set with $\alpha$ vertices, which is considered as a copy of $\Delta_{xy}$. The edge set $E_{H}:=\cup_{i=1}^{5}E_{i}$ is given by
$E_{1}=\{vw|v\in N_{x},w\in N_{y},vw\in E\},$
$E_{2}=\{vz^{\prime}_{i}|v\in N_{x},z^{\prime}_{i}\in\Delta^{\prime}_{xy},vz_{i}\in E\},$
$E_{3}=\{z_{i}w|z_{i}\in\Delta_{xy},w\in N_{y},z_{i}w\in E\},$
$E_{4}=\{z_{i}z^{\prime}_{j}|z_{i}z_{j}\in E,i\neq j,1\leq i,j\leq\alpha\},$
$E_{5}=\{e_{i}^{j}|e_{i}^{j}=z_{i}z^{\prime}_{i},1\leq i\leq\alpha,1\leq j\leq\beta-\alpha\},$
where $E_{5}$ contains $\beta-\alpha$ number of parallel edges between $z_{i}$ and $z^{\prime}_{i}$ for each $1\leq i\leq\alpha$. Notice that $\Delta_{xy}=\emptyset$ if $\alpha=0$ and $E_{5}=\emptyset$ if $\beta=\alpha$.
Lemma 3.2.
$H_{G}$ is $(\beta-1)$-regular.
Proof.
For any $v\in N_{x}$, we have $d(v,y)=2$ in $G$. Therefore, there are $\beta$ common neighbors of $v$ and $y$ in $V$, $\beta-1$ vertices of which lie in $N_{y}\cup\Delta_{xy}$. Thus, in the graph $H_{G}$, there are $\beta-1$ neighbors of $v$ in $N_{y}\cup\Delta^{\prime}_{xy}$ by definition. Similarly, for any $w\in N_{y}$, there are $\beta-1$ neighbors of $w$ in $N_{x}\cup\Delta_{xy}$.
For any $z_{i}\in\Delta_{xy}$, there are $\alpha$ common neighbors of $z_{i}$ and $y$ in $V$, $\alpha-1$ vertices of which lie in $N_{y}\cup\Delta_{xy}$. Together with $\beta-\alpha$ parallel edges between $z_{i}$ and $z^{\prime}_{i}$, there are $\beta-1$ neighbors of $z_{i}$ in $N_{y}\cup\Delta^{\prime}_{xy}$ in the graph $H_{G}$. Similarly, for any $z^{\prime}_{i}\in\Delta^{\prime}_{xy}$, there are $\beta-1$ neighbors of $z^{\prime}_{i}$ in $N_{x}\cup\Delta_{xy}$. ∎
Proof of Theorem 3.1.
By Theorem 2.6, $E_{H}$ can be decomposed into $\beta-1$ edge-disjoint perfect matchings. Since $|E_{5}|=\alpha(\beta-\alpha)$, there exists a perfect matching $\mathcal{M}$ such that
$|\mathcal{M}\cap E_{5}|\geq\left\lceil\frac{\alpha(\beta-\alpha)}{\beta-1}\right\rceil.$
We consider the following particular transport plan $\pi_{0}:V\times V\to[0,1]$ from $\mu_{x}^{\frac{1}{d+1}}$ to $\mu_{y}^{\frac{1}{d+1}}$:
$\pi_{0}(v,w)=\begin{cases}\frac{1}{d+1},&{\rm if}\ v\in N_{x}\cup\Delta_{xy},w\in N_{y}\ {\rm and}\ vw\in{\mathcal{M}};\\ \frac{1}{d+1},&{\rm if}\ v\in N_{x}\cup\Delta_{xy},w\in\Delta_{xy}\ {\rm and}\ vw^{\prime}\in{\mathcal{M}};\\ 0,&{\rm otherwise}.\end{cases}$
It is then direct to check that $\pi_{0}$ is indeed a transport plan. There are $|{\mathcal{M}}|$ pairs of $(v,w)$ such that $\pi_{0}(v,w)\neq 0$. Among them, there are $|{\mathcal{M}}\cap E_{5}|$ pairs with $d(v,w)=0$ and $|{\mathcal{M}}|-|{\mathcal{M}}\cap E_{5}|$ pairs with $d(v,w)=1$. Therefore, we have
$W\left(\mu_{x}^{\frac{1}{d+1}},\mu_{y}^{\frac{1}{d+1}}\right)$ $\leq\sum_{v\in V}\sum_{w\in V}d(v,w)\pi_{0}(v,w)$
$=\frac{1}{d+1}(|{\mathcal{M}}|-|{\mathcal{M}}\cap E_{5}|)$
$\leq\frac{1}{d+1}\left(d-1-\left\lceil\frac{\alpha(\beta-\alpha)}{\beta-1}\right\rceil\right).$
It follows by Theorem 2.3 that
$\kappa_{LLY}(x,y)=\frac{d+1}{d}\left(1-W\left(\mu_{x}^{\frac{1}{d+1}},\mu_{y}^{\frac{1}{d+1}}\right)\right)\geq\frac{2+\left\lceil\frac{\alpha(\beta-\alpha)}{\beta-1}\right\rceil}{d},$
completing the proof. ∎
Corollary 3.3.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $1\neq\beta\geq\alpha$. Then, the diameter of $G$ satisfies
$\mathrm{diam}(G)\leq\left\lfloor\frac{2d}{2+\left\lceil\frac{\beta-\alpha}{\beta-1}\alpha\right\rceil}\right\rfloor.$
Proof.
It is a direct consequence of Theorem 2.7 and Theorem 3.1. ∎
Remark 3.4.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. For any edge $xy\in E$, it has been proved that $\kappa_{LLY}(x,y)=\frac{2}{d}$ if $\beta\geq 2$ and $\alpha=0$ [24, Theorem 1.2], $\kappa_{LLY}(x,y)\geq\frac{2}{d}$ if $\beta=\alpha>1$ [24, Theorem 1.3], and $\kappa_{LLY}(x,y)\geq\frac{3}{d}$ if $\beta>\alpha\geq 1$ [22, Theorem 1.3]. Theorem 3.1 recovers and improves those previous results. In particular, for given $\alpha$ and $d$, our lower bound tends to the maximal possible curvature value $\frac{2+\alpha}{d}$ when $\beta$ is large enough.
Corollary 3.5.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta>\alpha^{2}-\alpha+1$. Then, for any edge $xy\in E$, we have $\kappa_{LLY}(x,y)=\frac{2+\alpha}{d}$, which is equivalent to say that there exists a perfect matching between $N_{x}$ and $N_{y}$.
Proof.
Since $\beta>\alpha^{2}-\alpha+1$, we have $\frac{\beta-\alpha}{\beta-1}\alpha>\alpha-1$. Hence, we derive from Theorem 3.1 that $\kappa_{LLY}(x,y)\geq\frac{2+\alpha}{d}$. By Theorem 2.4, we deduce that $\kappa_{LLY}(x,y)=\frac{2+\alpha}{d}$, or equivalently, there exists a perfect matching between $N_{x}$ and $N_{y}$. ∎
Remark 3.6.
For any edge $xy$ of an amply regular graph with parameters $(n,d,\alpha,\beta)$, several sufficient conditions involving restrictions on all parameters $n,d,\alpha,\beta$ for the existence of a perfect matching between $N_{x}$ and $N_{y}$ have been found in [12, Theorem 1.4]. In particular, a conjecture due to Bonini et. al. [8] on the existence of such local perfect matchings in conference graphs is confirmed. An advantage of Corollary 3.5, comparing to [12, Theorem 1.4], is that we only need a restriction on the two parameters $\alpha$ and $\beta$.
For an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $\alpha>\beta>1$, the Lin–Lu–Yau curvature is not necessarily positive anymore. For examples, two of the three locally Petersen graphs, Conway-Smith graph and Hall graph, are amply regular with parameters $(63,10,3,2)$ and $(65,10,3,2)$, respectively, and negative Lin–Lu–Yau curvature [13, Table 2]. We prove a positive lower curvature bound by assuming further $d\leq\alpha+\beta$ in the next subsection.
3.2 Case of $\alpha>\beta$ and $d\leq\alpha+\beta$
In this subsection, we prove the following estimate.
Theorem 3.7.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ with $\alpha>\beta$ and $d\leq\alpha+\beta$. Then the Lin–Lu–Yau curvature $\kappa_{LLY}(x,y)$ of any edge $xy\in E$ satisfies
$\kappa_{LLY}(x,y)\geq\frac{2}{d}.$
We construct a bipartite graph $H_{G}=(V_{H},E_{H})$ as follows. Let us denote the $\alpha$ vertices in $\Delta_{xy}$ by $z_{1},\cdots,z_{\alpha}$ and their copies by $\Delta^{\prime}_{xy}:=\{z^{\prime}_{1},\cdots,z^{\prime}_{\alpha}\}$. The vertex set of $H$ is given by
$V_{H}=N_{x}\cup N_{y}\cup\Delta_{xy}\cup\Delta^{\prime}_{xy}.$
The edge set $E_{H}:=\cup_{i=1}^{4}E_{i}$ is given by
$E_{1}=\{vw|v\in N_{x},w\in N_{y},vw\in E\},$
$E_{2}=\{vz^{\prime}_{i}|v\in N_{x},z^{\prime}_{i}\in\Delta^{\prime}_{xy},vz_{i}\in E\},$
$E_{3}=\{z_{i}w|z_{i}\in\Delta_{xy},w\in N_{y},z_{i}w\in E\},$
$E_{4}=\{z_{i}z^{\prime}_{j}|z_{i}z_{j}\in E,i\neq j,1\leq i,j\leq\alpha\}.$
Lemma 3.8.
In the graph $H_{G}$, the degree of each vertex in $N_{x}\cup N_{y}$ is $\beta-1$, and the degree of each vertex in $\Delta_{xy}\cup\Delta^{\prime}_{xy}$ is $\alpha-1$.
Proof.
For any $v\in N_{x}$, we have $d(v,y)=2$ in $G$. Therefore, there are $\beta$ common neighbors of $v$ and $y$ in $V$, $\beta-1$ vertices of which lie in $N_{y}\cup\Delta_{xy}$. Thus, in the graph $H_{G}$, there are $\beta-1$ neighbors of $v$ in $N_{y}\cup\Delta^{\prime}_{xy}$ by definition. Similarly, for any $w\in N_{y}$, there are $\beta-1$ neighbors of $w$ in $N_{x}\cup\Delta_{xy}$.
For any $z_{i}\in\Delta_{xy}$, there are $\alpha$ common neighbors of $z_{i}$ and $y$ in $V$, $\alpha-1$ vertices of which lie in $N_{y}\cup\Delta_{xy}$. Therefore, in the graph $H_{G}$, there are $\alpha-1$ neighbors of $z_{i}$ in $N_{y}\cup\Delta^{\prime}_{xy}$ by definition. Similarly, for any $z^{\prime}_{i}\in\Delta^{\prime}_{xy}$, there are $\alpha-1$ neighbors of $z^{\prime}_{i}$ in $N_{x}\cup\Delta_{xy}$. ∎
We recall the following lemma on existence of a perfect matching, which is a consequence of Hall’s marriage Theorem (Theorem 2.5). For the proof of the lemma, see, e.g., [12, Lemma 3.1].
Lemma 3.9.
Let $H=(V,E)$ be a bipartite graph with the bipartition $V=V_{1}\sqcup V_{2}$ such that $|V_{1}|=|V_{2}|$. Denote by $m:=|V_{1}|=|V_{2}|$. Suppose that
  • (i)
    $|\Gamma_{V_{2}}(S)|\geq|S|$ for any $S\subseteq V_{1}$ with $|S|\leq\frac{m+1}{2}$;
  • (ii)
    $|\Gamma_{V_{1}}(S)|\geq|S|$ for any $S\subseteq V_{2}$ with $|S|\leq\frac{m+1}{2}$.
Then, $H$ has a perfect matching.
We are now prepared to prove Theorem 3.7.
Proof of Theorem 3.7.
We claim that there is a perfect matching between $V_{1}:=N_{x}\cup\Delta_{xy}$ and $V_{2}:=N_{y}\cup\Delta^{\prime}_{xy}$. By Lemma 3.9 and by symmetry, we only need to prove that $|\Gamma_{V_{2}}(S)|\geq|S|$ for any $S\subseteq V_{1}$ with
$|S|\leq\frac{|V_{1}|+1}{2}=\frac{d}{2}.$
If $S\subseteq N_{x}$, it follows by Lemma 3.8 that
$|\Gamma_{V_{2}}(S)|\geq\beta-1\geq d-\alpha-1=|N_{x}|\geq|S|.$
If $S$ is not a subset of $N_{x}$, then $S$ contains at least one vertex in $\Delta_{xy}$. By Lemma 3.8, we have
$|\Gamma_{V_{2}}(S)|\geq\alpha-1\geq\left\lfloor\frac{\alpha+\beta}{2}\right\rfloor\geq\left\lfloor\frac{d}{2}\right\rfloor\geq|S|.$
Therefore, there is a perfect matching $\mathcal{M}$ between $V_{1}$ and $V_{2}$, as claimed. Following the definition of $\pi_{0}$ in the previous subsection, we have
$W\left(\mu_{x}^{\frac{1}{d+1}},\mu_{y}^{\frac{1}{d+1}}\right)\leq\sum_{v\in V}\sum_{w\in V}d(v,w)\pi_{0}(v,w)=\frac{1}{d+1}|{\mathcal{M}}|=\frac{d-1}{d+1}.$
Therefore, we derive by Theorem 2.3 that
$\kappa_{LLY}(x,y)=\frac{d+1}{d}\left(1-W\left(\mu_{x}^{\frac{1}{d+1}},\mu_{y}^{\frac{1}{d+1}}\right)\right)\geq\frac{2}{d},$
completing the proof. ∎
4 Bakry–Émery curvature formula of amply regular graphs
In this section, we calculate the Bakry–Émery curvature at a vertex of an amply regular graph associated with either a balanced or an anti-balanced signature. For that purpose, we first recall the general curvature matrix method developed in [16, 15, 20, 41, 42].
4.1 Curvature matrix
In [20], the second and third named authors establish the method for computing the Bakry–Émery curvature on connection graphs, i.e., graphs associated with an orthogonal or unitary matrix valued signature, building upon previous works [16, 15, 41, 42] on unsigned graphs. We recall the calculation procedure here. For our later purpose, we restrict ourselves to the current setting of a $d$-regular graph associated with a signature $\sigma:E\to\{+1,-1\}$. For any two vertices $x,y$, let $a_{xy}=1$ if $xy\in E$ and 00 otherwise.
Let $x\in V$ be a given vertex. For any integer $i$, define $S_{i}(x)$ to be the set of vertices at distance $i$ to $x$. Let us denote by
$S_{1}(x):=\{y_{1},y_{2},\ldots,y_{d}\},\,\text{and}\,\,S_{2}(x):=\{z_{1},z_{2},\ldots,z_{\ell}\}.$
For any $z_{k}\in S_{2}(x)$, we define its in-degree to be $d_{z_{k}}^{-}:=\sum_{i=1}^{d}a_{y_{i}z_{k}}$.
In the case of $N=\infty$, the curvature dimension inequality (2.1) at a vertex $x$ becomes
$\Gamma^{\sigma}_{2}(f)(x)-K\Gamma^{\sigma}(f)(x)\geq 0,$
for any function $f:V\rightarrow\mathbb{R}$. Observe that the value $\Gamma^{\sigma}_{2}(f)(x)$ depends only on the values of $f$ on $B_{2}(x)=\{x\}\cup S_{1}(x)\cup S_{2}(x)$. Let $\Gamma^{\sigma}_{2}(x)$ be the symmetric matrix of size $(1+d+|S_{2}(x)|)\times(1+d+|S_{2}(x)|)$ corresponding to the quadratic bilinear form $\Gamma^{\sigma}_{2}(f)(x)$.
Let us write $\Gamma^{\sigma}_{2}(x)$ blockwisely as
$\Gamma^{\sigma}_{2}(x)=\begin{pmatrix}\Gamma^{\sigma}_{2}(x)_{B_{1},B_{1}}&\Gamma^{\sigma}_{2}(x)_{B_{1},S_{2}}\\ \Gamma^{\sigma}_{2}(x)_{S_{2},B_{1}}&\Gamma^{\sigma}_{2}(x)_{S_{2},S_{2}}\end{pmatrix},$
where we use the notations that $B_{1}=B_{1}(x):=\{x\}\cup S_{1}(x)$ and $S_{2}=S_{2}(x)$. A key observation is that the lower right block
$\Gamma^{\sigma}_{2}(x)_{S_{2},S_{2}}=\frac{1}{4}\mathrm{diag}(d_{z_{1}}^{-},\ldots,d_{z_{\ell}}^{-}),$
is strictly positive definite. Let $Q(x)$ be the Schur complement of $\Gamma^{\sigma}_{2}(x)_{S_{2},S_{2}}$ in the matrix $\Gamma^{\sigma}_{2}(x)$, that is,
$Q(x):=\Gamma^{\sigma}_{2}(x)_{B_{1},B_{1}}-\Gamma^{\sigma}_{2}(x)_{B_{1},S_{2}}\left(\Gamma^{\sigma}_{2}(x)_{S_{2},S_{2}}\right)^{-1}\Gamma^{\sigma}_{2}(x)_{S_{2},B_{1}}.$
Below we recall the explicit formula of the $(d+1)\times(d+1)$ matrix $Q(x)$ from [20, (A.10)-(A.13)]: The rows and columns of $Q(x)$ are indexed by $x$ and $y_{1},\dots,y_{d}$. For convenience, we multiply $Q(x)$ by a factor $4$.
$4Q(x)_{x,x}=3d+d^{2}-\sum_{k=1}^{\ell}\frac{1}{d_{z_{k}}^{-}}\left(\sum_{i=1}^{d}a_{y_{i}z_{k}}\sigma_{xy_{i}}\sigma_{y_{i}z_{k}}\right)^{2},$ (4.1)
$4Q(x)_{x,y_{i}}=\sum_{j=1,j\neq i}^{d}a_{y_{j}y_{i}}\sigma_{xy_{j}}\sigma_{y_{j}y_{i}}-2\left(d+1\right)\sigma_{xy_{i}}$
$\hskip 156.49014pt+2\sum_{k=1}^{\ell}\frac{1}{d_{z_{k}}^{-}}\left(\sum_{j=1}^{d}a_{y_{j}z_{k}}\sigma_{xy_{j}}\sigma_{y_{j}z_{k}}\right)a_{y_{i}z_{k}}\sigma_{y_{i}z_{k}},\,\forall\,i,$ (4.2)
$4Q(x)_{y_{i},y_{i}}=\sum_{j=1,j\neq i}^{d}a_{y_{j}y_{i}}+2\left(d+1\right)-4\sum_{k=1}^{\ell}\frac{a_{y_{i}z_{k}}}{d_{z_{k}}^{-}},\,\forall\,i,$ (4.3)
$4Q(x)_{y_{i},y_{j}}=-4a_{y_{i}y_{j}}\sigma_{y_{i}y_{j}}+2\sigma_{xy_{i}}\sigma_{xy_{j}}-4\sum_{k=1}^{\ell}\frac{1}{d_{z_{k}}^{-}}a_{y_{i}z_{k}}a_{y_{j}z_{k}}\sigma_{y_{i}z_{k}}\sigma_{y_{j}z_{k}},\,\forall\,i,\,\forall\,j\neq i.$ (4.4)
We define another symmetric matrix $B$ of size $(d+1)\times(d+1)$ as below (see [20, (3.33)])
$B:=\begin{pmatrix}1&1&\cdots&1\\ &1&&\\ &&\ddots&\\ &&&1\end{pmatrix}$ (4.5)
We then calculate the product $2BQ(x)B^{\top}$ and write it blockwisely as
$2BQ(x)B^{\top}=\begin{pmatrix}a&\omega^{\top}\\ \omega&(2BQ(x)B^{\top})_{\hat{1}}\end{pmatrix},$
where $(2BQ(x)B^{\top})_{\hat{1}}$ stands for the $d\times d$ matrix obtained from $2BQ(x)B^{\top}$ by removing the first row and the first column. By [20, Proposition A.1], it always holds that $a\geq 0$. Taking the Schur complement of $a$ in the matrix $2BQ(x)B^{\top}$, we arrive at the so-called curvature matrix [20, Definition 3.1].
Definition 4.1 (Curvature matrix).
Let $(G,\sigma)$ be a signed graph with $G=(V,E)$ being $d$-regular. The curvature matrix at a vertex $x\in V$ is defined as
$A_{\infty}(G,\sigma,x):=(2BQ(x)B^{\top})_{\hat{1}}-\omega a^{\dagger}\omega^{\top},$
where $a^{\dagger}=\frac{1}{a}$ if $a>0$ and 00 if $a=0$.
The Bakry–Émery curvature at a vertex of a signed graph is actually equal to the smallest eigenvalue of the curvature matrix.
Theorem 4.2 ([20, Theorem 3.2]).
Let $(G,\sigma)$ be a signed graph with $G=(V,E)$ being $d$-regular. Then we have
$K_{BE}(G,\sigma,x,\infty)=\lambda_{\min}(A_{\infty}(G,\sigma,x)),$
where $\lambda_{\min}(\cdot)$ stands for the smallest eigenvalue of the matrix.
4.2 Bakry–Émery curvature formula with balanced signatures
In this subsection, we derive the following formula for the Bakry–Émery curvature of an amply regular graph $G=(V,E)$. Note that $G$ can be considered as a signed graph $(G,\sigma)$ with $\sigma$ being the all-$+1$ signature.
Theorem 4.3.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Let $x\in V$ be a vertex, and $A_{S_{1}(x)}$ be the adjacency matrix of the local graph induced by $S_{1}(x)$. Then we have
$K_{BE}(+,x)=2+\frac{\alpha}{2}+\left(\frac{2d(\beta-2)-\alpha^{2}}{2\beta}+\frac{2}{\beta}\min_{\lambda\in\mathrm{sp}\left(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}}\right)}\left(\lambda-\frac{\alpha}{2}\right)^{2}\right)_{-},$ (4.6)
where the minimum is taken over every eigenvalue of $A_{S_{1}(x)}$ with an eigenvector vertical to $\mathbf{1}$, and $a_{-}:=\min\{0,a\}$, for any $a\in\mathbb{R}$.
Remark 4.4.
Let $\mathrm{sp}\left(A_{S_{1}(x)}\right):=\{\lambda_{1},\cdots,\lambda_{d-1},\lambda_{d}\}$ be the multiset of adjacency eigenvalues counting multiplicity. Then the set $\mathrm{sp}\left(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}}\right)=\mathrm{sp}\left(A_{S_{1}(x)}\right)\setminus\{\alpha\}$. Notice that, if the local graph induced by $S_{1}(x)$ is disconnected, then the set $\mathrm{sp}\left(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}}\right)$ still contains the eigenvalue $\alpha$.
Proof.
For $x\in V$, recall our notations $S_{1}(x)=\{y_{1},\ldots,y_{d}\}$ and $S_{2}(x)=\{z_{1},\ldots,z_{\ell}\}$. Taking the signature $\sigma$ to be the all-$+1$ signature, we derive the matrix $4Q(x)$ from (4.1)-(4.4) as follows.
$4Q(x)=\begin{pmatrix}(4+\alpha)d&-4-\alpha&\cdots&-4-\alpha\\ -4-\alpha&&&\\ \vdots&&P(x)&\\ -4-\alpha&&&\end{pmatrix},$
where the $d\times d$ submatrix $P(x)$ is given by
$(P(x))_{ii}=$ $2d+\alpha+2-\frac{4}{\beta}(d-\alpha-1),\,\,\,i=1,2,\ldots,d,$ (4.7)
$(P(x))_{ij}=$ $2-4a_{y_{i}y_{j}}-\frac{4}{\beta}\sum_{k=1}^{\ell}a_{y_{i}z_{k}}a_{z_{k}y_{j}},\,\,\,i\neq j.$ (4.8)
In the above, we have applied the fact that $d_{z_{k}}^{-}=\beta$ and
$\sum_{k=1}^{\ell}d_{z_{k}}^{-}=\sum_{i=1}^{d}(d-\alpha-1)=d(d-\alpha-1).$
Recalling the definition of matrix $B$ in (4.5), we calculate the product
$4BQ(x)B^{\top}=\begin{pmatrix}0&0&\cdots&0\\ 0&&&\\ \vdots&&P(x)\\ 0&&&\end{pmatrix}.$
By Definition 4.1, the matrix $\frac{1}{2}P(x)$ is the curvature matrix at $x$. Hence, due to Theorem 4.2, we have
$K_{BE}(+,x)=\frac{1}{2}\lambda_{\min}(P(x)).$ (4.9)
We observe that for $i\neq j$
$\sum_{w\in V}a_{y_{i}w}a_{wy_{j}}=1+\left(A^{2}_{S_{1}(x)}\right)_{ij}+\sum_{k=1}^{\ell}a_{y_{i}z_{k}}a_{z_{k}y_{j}}=\left\{\begin{array}[]{ll}\alpha,&\hbox{if $a_{y_{i}y_{j}}=1$;}\\ \beta,&\hbox{if $a_{y_{i}y_{j}}=0$.}\end{array}\right.$
That is, we have
$\sum_{k=1}^{\ell}a_{y_{i}z_{k}}a_{z_{k}y_{j}}=\alpha\left(A_{S_{1}(x)}\right)_{ij}+\beta\left(1-\left(A_{S_{1}(x)}\right)_{ij}\right)-1-\left(A^{2}_{S_{1}(x)}\right)_{ij}.$ (4.10)
Inserting (4.10) into (4.8) yields for $i\neq j$ that
$(P(x))_{ij}=-\frac{2(\beta-2)}{\beta}-\frac{4\alpha}{\beta}\left(A_{S_{1}(x)}\right)_{ij}+\frac{4}{\beta}\left(A^{2}_{S_{1}(x)}\right)_{ij}.$ (4.11)
Let us rearrange (4.7) accordingly as
$(P(x))_{ii}=\frac{2(\beta-2)(d-1)}{\beta}+\frac{4\alpha}{\beta}+(4+\alpha).$ (4.12)
Combining (4.11) and (4.12), we obtain the curvature matrix
$\frac{1}{2}P$ $=\left(\frac{\beta-2}{\beta}(d-1)+\frac{2\alpha}{\beta}+2+\frac{\alpha}{2}\right)\mathbf{I}_{d}+\frac{\beta-2}{\beta}(\mathbf{I}_{d}-\mathbf{J}_{d})-\frac{2\alpha}{\beta}A_{S_{1}(x)}+\frac{2}{\beta}\left(A_{S_{1}(x)}^{2}-\alpha\mathbf{I}_{d}\right)$
$=\left(\frac{d(\beta-2)}{\beta}+2+\frac{\alpha}{2}\right)\mathbf{I}_{d}-\frac{\beta-2}{\beta}\mathbf{J}_{d}-\frac{2\alpha}{\beta}A_{S_{1}(x)}+\frac{2}{\beta}A_{S_{1}(x)}^{2},$
where $\mathbf{I}_{d}$ and $\mathbf{J}_{d}$ are the $d\times d$ identity matrix and all-$1$ matrix, respectively, and we used the fact $\left(A_{S_{1}(x)}^{2}\right)_{ii}=\alpha$.
Since the local graph induced by $S_{1}(x)$ is $\alpha$-regular, the constant vector $\mathbf{1}$ is an eigenvector of its adjacency matrix $A_{S_{1}(x)}$ with eigenvalue $\alpha$. It is then straightforward to check that
$\frac{1}{2}P(x)\mathbf{1}=\left(2+\frac{\alpha}{2}\right)\mathbf{1}.$
The eigenvalues of $\frac{1}{2}P(x)$ corresponding eigenvectors vertical to the constant vector $\mathbf{1}$ are given by
$2+\frac{\alpha}{2}+\frac{d(\beta-2)}{\beta}-\frac{2\alpha}{\beta}\lambda+\frac{2}{\beta}\lambda^{2}=2+\frac{\alpha}{2}+\frac{d(\beta-2)-\alpha^{2}}{\beta}-\frac{2}{\beta}\left(\lambda-\frac{\alpha}{2}\right)^{2},$ (4.13)
where $\lambda\in\mathrm{sp}(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}})$ is any eigenvalue of $A_{S_{1}(x)}$ corresponding to an eigenvector vertical to $\mathbf{1}$. By (4.9), we arrive at
$K_{BE}(+,x)=2+\frac{\alpha}{2}+\left(\frac{2d(\beta-2)-\alpha^{2}}{2\beta}+\frac{2}{\beta}\min_{\lambda\in\mathrm{sp}\left({A_{S_{1}(x)}|}_{\mathbf{1}^{\perp}}\right)}\left(\lambda-\frac{\alpha}{2}\right)^{2}\right)_{-}.$
This completes the proof. ∎
Consequently, we obtain the Bakry–Émery curvature of an amply regular graph with parameters $(n,d,\alpha,\beta)$, under various restriction on the relation between $\alpha$ and $\beta$.
Corollary 4.5.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Let $x$ be a vertex.
  • (i)
    If $\beta=1$, then $K_{BE}(+,x)=2+\frac{\alpha}{2}-d$.
  • (ii)
    If $1\neq\beta\geq\alpha$ and $(\alpha,\beta)\neq(2,2)$, then $K_{BE}(+,x)=2+\frac{\alpha}{2}$.
  • (iii)
    If $\alpha=\beta=2$, then $2\leq K_{BE}(+,x)\leq 3$. Moreover, the upper bound is achieved if and only if the local graph induced by $S_{1}(x)$ is a disjoint union of at least two $3$-cycles, and the lower bound is achieved if and only if the local graph contains a $m$-cycle with $m$ being a positive multiple of $6$.
Remark 4.6.
A bit surprisingly, the Bakry–Émery curvature of any amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$ is uniquely determined by its parameters, except for only one case $(\alpha,\beta)=(2,2)$. The $4\times 4$ Rooks’ graph and the Shrikhande graph are both strongly regular with parameters $(16,6,2,2)$, i.e., $(\alpha,\beta)=(2,2)$. It has been noticed in the arXiv version of [16, Example 11.4] that the local graph of any vertex $x$ in the $4\times 4$ Rooks’ graph is a disjoint union of two $3$-cycles and $K_{BE}(+,x)=3$, while the local graph of any vertex $x$ in the Shrikhande graph is a $6$-cycle and $K_{BE}(+,x)=2$.
Proof.
Suppose $\beta=1$. Then local graph induced by $S_{1}(x)$ is a disjoint union of complete graphs $K_{\alpha+1}$ of size $\alpha+1$ with at least two connected components. Therefore, the set of adjacency eigenvalues $\mathrm{sp}(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}})$ consists of $\alpha$ and $-1$. Hence, we have
$\min_{\lambda\in\mathrm{sp}\left({A_{S_{1}(x)}|}_{\mathbf{1}^{\perp}}\right)}\left(\lambda-\frac{\alpha}{2}\right)^{2}=\frac{\alpha^{2}}{4}.$
Therefore, we derive from (4.6) that $K_{BE}(+,x)=2+\frac{\alpha}{2}-d$ if $\beta=1$. This proves $(i)$.
Next, we show $(ii)$. Notice that $d\geq\alpha+1$. Indeed, if $d=\alpha+1$, the graph has to be complete. Therefore, we have $d\geq\alpha+2$. If $\beta\geq\alpha\geq 3$, then we estimate
$2d(\beta-2)-\alpha^{2}\geq 2(\alpha+2)(\alpha-2)-\alpha^{2}=\alpha^{2}-8\geq 0$
Therefore, we derive from the formula (4.6) that $K_{BE}(+,x)=2+\frac{\alpha}{2}$ if $\beta\geq\alpha\geq 3$.
If $\alpha=0$ and $\beta\geq 2$, we have $2d(\beta-2)-\alpha^{2}=2d(\beta-2)\geq 0$ and hence $K_{BE}(+,x)=2$.
If $\alpha=1$ and $\beta\geq 2$, then the local graph induced by $S_{1}(x)$ is a $1$-regular graph, and hence must be a disjoint union of complete graphs $K_{2}$ of size $2$. If the local graph has only one connected component, then the graph $G$ is the complete graph $K_{3}$. Therefore, the local graph has at least two connected components. This leads to
$\min_{\lambda\in\mathrm{sp}\left({A_{S_{1}(x)}|}_{\mathbf{1}^{\perp}}\right)}\left(\lambda-\frac{\alpha}{2}\right)^{2}=\frac{1}{4}.$
By (4.6), we obtain $K_{BE}(+,x)=2+\frac{\alpha}{2}$ if $\alpha=1$ and $\beta\geq 2$.
If $\alpha=2$ and $\beta\geq 3$, we have $02d(\beta-2)-\alpha^{2}\geq 2d-\alpha^{2}>$ 0, and hence $K_{BE}(+,x)=2+\frac{\alpha}{2}$. This completes the proof of $(ii)$.
Finally, let us consider the case $\alpha=\beta=2$. Then the formula (4.6) tells that
$K_{BE}(+,x)=3+\left(-1+\min_{\lambda\in\mathrm{sp}\left({A_{S_{1}(x)}}|_{\mathbf{1}^{\perp}}\right)}\left(\lambda-1\right)^{2}\right)_{-}.$
It follows that $2\leq K_{BE}(+,x)\leq 3$, where $K_{BE}(+,x)=3$ holds if and only if
$\min_{\lambda\in\mathrm{sp}\left({A_{S_{1}(x)}}|_{\mathbf{1}^{\perp}}\right)}\left(\lambda-1\right)^{2}\geq 1,$ (4.14)
and $K_{BE}(+,x)=2$ holds if and only if
$\min_{\lambda\in\mathrm{sp}\left({A_{S_{1}(x)}}|_{\mathbf{1}^{\perp}}\right)}\left(\lambda-1\right)^{2}=0.$ (4.15)
Notice that the local graph induced by $S_{1}(x)$ is a $2$-regular graph, and hence a disjoint union of cycles. We first observe that the local graph can not contain a $4$-cycle, since otherwise there exist two vertices at distance $2$ from this $4$-cycle having at least $3$ common neighbors, contradicting the assumption that $\beta=2$.
Recall that the eigenvalues of a cycle $C_{n}$ of $n$ vertices are $2\cos\frac{2j\pi}{n},j=0,1,\ldots,\left\lfloor\frac{n}{2}\right\rfloor$. We have $K_{BE}(+,x)=3$, or equivalently (4.14) holds, if and only if the local graph is a disjoint union of at least two $3$-cycles.
We have $K_{BE}(+,x)=2$, or equivalently (4.15) holds, if and only if the local graph has an eigenvalue $1$, which can only happen when the local graph contains a $m$-cycle such that $m$ is a positive multiple of $6$. ∎
Corollary 4.7.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Then, the diameter of $G$ satisfies
  • (i)
    If $2(\beta-2)d\geq\alpha^{2}$, then
    $\mathrm{diam}(G)\leq\left\lfloor\frac{2d}{2+\frac{\alpha}{2}}\right\rfloor.$ (4.16)
  • (ii)
    If $2(\beta-2)d\geq\alpha^{2}-\alpha\beta$, then
    $\mathrm{diam}(G)\leq d.$ (4.17)
Proof.
(i) If $2(\beta-2)d\geq\alpha^{2}$, it follows by Theorem 4.3 that $K_{BE}(+,x)=2+\frac{\alpha}{2}$ for any vertex $x\in V$. Then, Theorem 2.11 yields inequality (4.16).
(ii) If $2(\beta-2)d\geq\alpha^{2}-\alpha\beta$, it follows by Theorem 4.3 that $K_{BE}(+,x)\geq 2$ for any vertex $x\in V$. Then, Theorem 2.11 yields inequality (4.17). ∎
Now we are prepared to show Theorem 1.3, Corollary 1.4 and Theorem 1.5 from the Introduction.
Proof of Theorem 1.3.
By Corollary 4.5 and Theorem 2.11, we have
$\mathrm{diam}(G)\leq\left\lfloor\frac{2d}{2+\frac{\alpha}{2}}\right\rfloor.$
Combining the above estimate with Theorem 3.3, we obtain the desired conclusion. ∎
Proof of Corollary 1.4.
This is a direct consequence of Theorem 1.3. We mention that in the case $2=\beta=\alpha>\varepsilon d\geq 1$, we have by Corollary 4.7 that $\mathrm{diam}(G)\leq d<2/\varepsilon<4/\varepsilon$. ∎
Proof of Theorem 1.5.
There are finitely many amply regular graphs with diameter $2$ and fixed degree. We suppose that $G$ has diameter at least three. If $d<3(\alpha-\beta+2)$, it follows by Corollary 2.15 and Theorem 2.14 that $G$ is the icosahedron or the line graph of a regular graph of girth at least $5$. If $G$ is the line graph of a regular graph of girth at least $5$, then $\beta=1$, which is contradictory to the assumption that $\alpha\leq 6\beta-9$.
Now, we suppose that $d\geq 3(\alpha-\beta+2)$. If $\alpha\geq\beta$, since $6\beta-9-\alpha\geq 0$, we have
$\frac{2d(\beta-2)-\alpha^{2}}{2\beta}\geq\frac{6(\alpha-\beta+2)(\beta-2)-\alpha^{2}}{2\beta}$
$=$ $\frac{(6\beta-9-\alpha)(\alpha-\beta+3)+3-(\alpha+3)\beta}{2\beta}>-\frac{\alpha+3}{2}.$
Then, Theorem 4.3 shows that $K_{BE}(+,x)>\frac{1}{2}$. Combining with Corollary 4.5, we have $K_{BE}(+,x)>\frac{1}{2}$ whenever $a\leq 6\beta-9$. The positive curvature lower bound implies, by Theorem 2.11, that $G$ is finite and has diameter at most $4d$. This completes the proof. ∎
4.3 Bakry–Émery curvature formula with anti-balanced signatures
In this subsection, we study the Bakry–Émery curvature of an amply regular graph associated with the all-$-1$ signature. By Proposition 2.10, this is equivalent to consider any anti-balanced signature on an amply regular graph.
Theorem 4.8.
Let $(G,-)$ be a signed graph, where $G=(V,E)$ is an amply regular graph with parameters $(n,d,\alpha,\beta)$, and -- stands for the all-$-1$ signature. Let $x\in V$ be a vertex, and $A_{S_{1}(x)}$ be the adjacency matrix of the local graph induced by $S_{1}(x)$. Then we have
$K_{BE}(-,x)=2+\left(\frac{2d(\beta-2)-\alpha^{2}}{2\beta}+\frac{5}{2}\alpha-2\beta+\frac{2}{\beta}\min_{\lambda\in\mathrm{sp}\left(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}}\right)}\left(\lambda+\beta-\frac{\alpha}{2}\right)^{2}\right)_{-},$ (4.18)
where the minimum is taken over every eigenvalue of $A_{S_{1}(x)}$ with an eigenvector vertical to $\mathbf{1}$, and $a_{-}:=\min\{0,a\}$, for any $a\in\mathbb{R}$.
Proof.
We first switch the all-$-1$ signature to be an anti-balanced signature $\sigma$ such that
$\sigma_{xy_{i}}$ $=+1,\,\,\text{for any $y_{i}\in S_{1}(x)$},$
$\sigma_{y_{i}y_{j}}$ $=-1,\,\,\text{for any edge $y_{i}y_{j}$, where $y_{i},y_{j}\in S_{1}(x)$},$
$\sigma_{y_{i}z_{k}}$ $=+1,\,\,\text{for any edge $y_{i}z_{k}$, where $y_{i}\in S_{1}(x)$ and $z_{k}\in S_{2}(x)$}.$
This can be achieved by choosing a switching function $\tau:V\to\{+1,-1\}$ such that $\tau(x)=+1,\tau|_{S_{1}(x)}\equiv-1$ and $\tau|_{S_{2}(x)}\equiv+1$. By Proposition 2.10, we have
$K_{BE}(-,x)=K_{BE}(G,\sigma,x,\infty).$ (4.19)
Next we apply Theorem 4.2 to calculate $K_{BE}(G,\sigma,x,\infty)$. First, we derive from (4.1)-(4.4) the $(d+1)\times(d+1)$ matrix $4Q(x)$ as follows:
$4Q(x)=\begin{pmatrix}(4+\alpha)d&-4-3\alpha&\cdots&-4-3\alpha\\ -4-3\alpha&&&\\ \vdots&&R(x)&\\ -4-3\alpha&&&\end{pmatrix},$
where $R(x)$ is a $d\times d$ submatrix given by
$(R(x))_{ii}$ $=2d+\alpha+2-\frac{4}{\beta}(d-\alpha-1),\,\,\,i=1,\dots,d.$
$(R(x))_{ij}$ $=2+4a_{y_{i}y_{j}}-\frac{4}{\beta}\sum_{k=1}^{\ell}a_{y_{i}z_{k}}a_{z_{k}y_{j}},\,\,\,i\neq j.$
Recall the definition of matrix $B$, we calculate the product
$4BQ(x)B^{\top}=\begin{pmatrix}4\alpha d&6\alpha&\cdots&6\alpha\\ 6\alpha&&&&\\ \vdots&&R(x)&&\\ 6\alpha&&&&\\ \end{pmatrix}$
By Definition 4.1, we obtain the curvature matrix
$2A_{\infty}(G,-,x)$ $=R(x)-\begin{pmatrix}6\alpha\\ 6\alpha\\ \vdots\\ 6\alpha\end{pmatrix}\frac{1}{4\alpha d}\begin{pmatrix}6\alpha&6\alpha&\cdots&6\alpha\end{pmatrix}$
$=R(x)-\frac{9\alpha}{d}\mathbf{J}_{d},$
where $\mathbf{J}_{d}$ is the $d\times d$ all-$1$ matrix. In the remaining part of the proof, we will write $A_{\infty}(x):=A_{\infty}(G,-,x)$ for short.
Applying (4.10), we derive that for $i\neq j$
$(R(x))_{ij}=-2+8\left(A_{S_{1}(x)}\right)_{ij}-\frac{4\alpha}{\beta}\left(A_{S_{1}(x)}\right)_{ij}+\frac{4}{\beta}+\frac{4}{\beta}\left(A_{S_{1}(x)}^{2}\right)_{ij}.$
Therefore, the curvature matrix is
$A_{\infty}(x)=\frac{2}{\beta}A_{S_{1}(x)}^{2}+\left(4-\frac{2\alpha}{\beta}\right)A_{S_{1}(x)}+\left(2+\frac{\alpha}{2}+d-\frac{2d}{\beta}\right)\mathbf{I}_{d}+\left(\frac{2}{\beta}-\frac{9\alpha}{2d}-1\right)\mathbf{J}_{d}.$
Then, we check for the constant vector $\mathbf{1}$ that
$A_{\infty}(x)\mathbf{1}=2\mathbf{1}.$
All the other eigenvalue of $A_{\infty}(x)$ corresponding to an eigenvector vertical to $\mathbf{1}$ is given by
$\frac{2}{\beta}\lambda^{2}+\left(4-\frac{2\alpha}{\beta}\right)\lambda+2+\frac{\alpha}{2}+d-\frac{2d}{\beta}$
$=$ $2+\frac{2d(\beta-2)-\alpha^{2}}{2\beta}+\frac{5\alpha}{2}-2\beta+\frac{2}{\beta}\left(\lambda+\beta-\frac{\alpha}{2}\right)^{2},$
where $\lambda\in\mathrm{sp}(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}})$ is an eigenvalue of $A_{S_{1}(x)}$ corresponding to an eigenvector vertical to $\mathbf{1}$. In conclusion, we have the Bakry–Émery curvature by (4.19) and Theorem 4.2
$K_{BE}(-,x)=2+\left(\frac{2d(\beta-2)-\alpha^{2}}{2\beta}+\frac{5\alpha}{2}-2\beta+\frac{2}{\beta}\min_{\lambda\in\mathrm{sp}\left(A_{S_{1}(x)}|_{\mathbf{1}^{\top}}\right)}\left(\lambda+\beta-\frac{\alpha}{2}\right)^{2}\right)_{-}.$
This completes the proof. ∎
Corollary 4.9.
For an amply regular graph $G=(V,E)$ with parameters $(n,d,\alpha,1)$ associated an anti-balanced signature, the Bakry–Émery curvature at a vertex $x\in V$ satisfies
$K_{BE}(-,x)=2+\left(-d-2+\frac{5\alpha}{2}\right)_{-}.$
Proof.
Since $\beta=1$, the local graph induced by $S_{1}(x)$ is a disjoint union of complete graphs $K_{\alpha+1}$, with at least $2$ connected components. Therefore, we have
$\min_{\lambda\in\mathrm{sp}\left(A_{S_{1}(x)}|_{\mathbf{1}^{\perp}}\right)}\left(\lambda+1-\frac{\alpha}{2}\right)^{2}=\frac{\alpha^{2}}{2}.$
Applying Theorem 4.8 yields
$K_{BE}(-,x)=2+\left(-d-2+\frac{5}{2}\alpha\right)_{-}.$
This completes the proof. ∎
5 Applications in bounding eigenvalues, isoperimetric constants and volume growth of amply regular graphs
In this section, we discuss further applications of our discrete Ricci curvature estimates.
5.1 Application in bounding the eigenvalues
In this subsection, we consider finite amply regular graphs. In fact, the finiteness will be guaranteed by the parameter conditions we impose.
Theorem 5.1.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Let
$-d\leq\theta_{1}\leq\cdots\leq\theta_{n-1}\leq\theta_{n}=d$
be the eigenvalues of its adjacency matrix.
  • (i)
    If $2d(\beta-2)\geq\alpha^{2}$, then
    $\theta_{n-1}\leq d-2-\frac{\alpha}{2}.$
  • (ii)
    If $2d(\beta-2)\geq\alpha(\alpha-\beta)$, then
    $\theta_{n-1}\leq d-2.$
  • (iii)
    If $2d(\beta-2)\geq(\alpha-\beta)(\alpha-4\beta)$, then
    $\theta_{1}\geq-d+2.$
The Bakry–Émery curvature of a signed graph $(G,\sigma)$ is closely related to eigenvalue estimates for the Laplacian $\Delta^{\sigma}$. We say $\lambda^{\sigma}\in\mathbb{R}$ an eigenvalue of $\Delta^{\sigma}$ if there exists a non-zero function $f:V\to\mathbb{R}$ such that $\Delta^{\sigma}f+\lambda^{\sigma}f=0$. As a preparation for proving Theorem 5.1, we recall the following Lichnerowicz type eigenvalue estimate.
Theorem 5.2 ([29, Theorem 1.9]).
Assume a finite signed graph $(G,\sigma)$ satisfies $CD^{\sigma}(K,\infty)$ for $K\in\mathbb{R}$. Then we have for any non-zero eigenvalue $\lambda^{\sigma}$ of $\Delta^{\sigma}$ that $\lambda^{\sigma}\geq K$.
Recalling $\Delta^{\sigma}=A^{\sigma}-D$, the eigenvalues of the matrix $-\Delta^{\sigma}$ can be easily translated into eigenvalues of the signed adjacency matrix. We mention that the smallest eigenvalue of $\Delta^{\sigma}$ is 00 if and only if $(G,\sigma)$ is balanced. Therefore, we have the following observations.
Lemma 5.3.
Let $(G,\sigma)$ be a finite signed graph with $G$ being $d$-regular. Let $\lambda^{\sigma}$ be the smallest non-zero eigenvalue of $\Delta^{\sigma}$ and $\theta_{1}\leq\cdots\leq\theta_{n-1}\leq\theta_{n}=d$ be the eigenvalues of adjacency matrix of the underlying graph $G$.
  • (i)
    If $\sigma$ is the all-$+1$ signature, then $\lambda^{\sigma}=d-\theta_{n-1}$.
  • (ii)
    If $\sigma$ is the all-$-1$ signature, then $\lambda^{\sigma}=d+\theta_{1}$.
The proof of the above lemma is straightforward, which we omit.
Proof of Theorem 5.1.
(i) Recall Theorem 4.3. If $2d(\beta-2)-\alpha^{2}\geq 0$, then we obtain $K_{BE}(+,x)=2+\frac{\alpha}{2}.$ Therefore, we derive by applying Theorem 5.2 and Lemma 5.3 in the case of $\sigma\equiv+1$ that
$\theta_{n-1}=d-\lambda^{\sigma}\leq d-2-\frac{\alpha}{2}.$
(ii) If $2d(\beta-2)\geq\alpha^{2}-\alpha\beta$, then we have $\frac{2d(\beta-2)}{2\beta}\geq-\frac{\alpha}{2}$ and hence $K_{BE}(+,x)\geq 2$. Therefore, we derive by Theorem 5.2 and Lemma 5.3 that
$\theta_{n-1}=d-\lambda^{\sigma}\leq d-2.$
(iii) Recall Theorem 4.8. If $2d(\beta-2)\geq(\alpha-\beta)(\alpha-4\beta)$, then
$\frac{2d(\beta-2)-\alpha^{2}}{2\beta}+\frac{5\alpha}{2}-2\beta\geq 0,$
and therefore, $K_{BE}(-,x)=2$. Then, we derive by applying Theorem 5.2 and Lemma 5.3 in the case of $\sigma\equiv-1$ that
$\theta_{1}=-d+\lambda^{\sigma}\geq-d+2.$
We complete the proof. ∎
Next, we prove Theorem 1.7 from the Introduction.
Proof of Theorem 1.7.
By Theorem 5.1(iii), we only need to prove that
$(\alpha-\beta)(\alpha-4\beta)\leq 2d(\beta-2).$
One one hand, suppose that $2\leq\alpha\leq\beta$. By Theorem 2.16, we have $d\geq 2\beta$. An elementary computation shows that
$(\alpha-\beta)(\alpha-4\beta)\leq 4\beta(\beta-2)\leq 2d(\beta-2).$
On the other hand, suppose that $\beta<\alpha\leq 10\beta-12$. If $d<3(\alpha-\beta+2)$, Corollary 2.15 and Theorem 2.14 implies that $G$ is the icosahedron or the line graph of a regular graph of girth at least $5$. Since the diameter of $G$ is at least $4$, $G$ is not the icosahedron. If $G$ is the line graph of a regular graph of girth at least $5$, then $\beta=1$, which is contradictory to the assumption that $\alpha\leq 10\beta-12$. Thus, we have $d\geq 3(\alpha-\beta+2)$. Then
$(\alpha-\beta)(\alpha-4\beta)-2d(\beta-2)$ $\leq(\alpha-\beta)(\alpha-4\beta)-6(\alpha-\beta+2)(\beta-2)$
$=(\alpha-\beta)(\alpha-10\beta+12)-12\beta+24\leq 0.$
We complete the proof. ∎
Besides, a positive Lin–Lu–Yau curvature lower bound also gives a Licnherowicz type eigenvalue estimate. We state below the result restricting to the case of a regular graph.
Theorem 5.4 ([25, Theorem 4.2]).
Let $G=(V,E)$ be a $d$-regular finite graph and $\lambda$ be a nonzero eigenvalue of the Laplacian $\Delta$. If $\kappa_{LLY}(x,y)\geq k>0$ for any edge $xy\in E$, then
$\lambda\geq dk.$
Combining both estimates derived from the Bakry–Émery curvature and Lin–Lu–Yau curvature bounds on amply regular graphs, we obtain Theorem 1.6.
Proof of Theorem 1.6.
(i) It follows by Theorem 3.1 and Theorem 5.4 that
$\theta_{n-1}\leq d-2-\left\lceil\frac{(\beta-\alpha)\alpha}{\beta-1}\right\rceil.$
It follows by Corollary 4.5, Theorem 5.2 and Lemma 5.3 that
$\theta_{n-1}\leq d-2-\frac{\alpha}{2}.$
Combining these two estimates, we obtain the desired result.
(ii) The result follows directly by Corollary 4.5, Theorem 5.2 and Lemma 5.3. ∎
5.2 Application in bounding the isoperimetric constants
We first give the definitions of two important isoperimetric constants on graphs. For any two vertex set $A$ and $B$ of a graph $G=(V,E)$, let us denote by ||$\cdot$|| ⋅ | the cardinality of a set, and by $|E(A,B)|:=\sum_{u\in A}\sum_{v\in B}a_{uv}$, where $a_{uv}=1$ if $uv\in E$ and 00 otherwise. Notice that $|E(A,A)|$ is twice the number of edges with both ends in $A$.
Definition 5.5 (Edge Cheeger constant).
Let $G=(V,E)$ be a $d$-regular connected graph, the edge Cheeger constant $h$ is defined as
$h:=\inf_{\Omega\subset V,|\Omega|\leq\frac{1}{2}|V|}\frac{|E(\Omega,\Omega^{c})|}{d|\Omega|}.$
This constant is also called edge expansion, measuring how fast this graph expand.
Definition 5.6 ([6, 47], Bipartiteness constant).
Let $G=(V,E)$ be a $d$-regular connected graph, the bipartiteness constant $\beta_{0}$ is defined as
$\beta_{0}:=\inf_{L,R\subset V,L\cap R\neq\emptyset}\frac{|E(L,L)|+|E(R,R)|+|\partial(L\cap R)|}{d|L\cup R|}.$
Here the notation $\partial(\Omega):=E(\Omega,\Omega^{c})$ is the edge boundary set of a subset $\Omega$.
Notice that $\beta_{0}=0$ if and only if the graph $G$ is bipartite. The constant $\beta_{0}$ measures how far a graph is from being bipartite. The bipartite constant is introduced independently by Bauer and Jost [6] and Trevisan [47]. Bauer and Jost call the constant $1-\beta_{0}$ dual Cheeger constant.
Both $h$ and $\beta_{0}$ are closed related to eigenvalues of Laplacians via so-called Cheeger and dual Cheeger inequalities, for which we refer to [6] and references therein. Translating into adjacency eigenvalues, there holds $dh^{2}/2\leq d-\theta_{n-1}\leq 2dh,\,\,\text{and}\,\,d\beta_{0}^{2}/2\leq d+\theta_{1}\leq 2d\beta_{0}$.
Recalling the eigenvalue estimates on amply regular graphs from above, we further derive estimates for edge Cheeger and bipartiteness constants.
Corollary 5.7.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$.
  • (i)
    Suppose $1\neq\beta\geq\alpha$. Then for any subset $\Omega\subseteq V$ with $|\Omega|\leq\frac{1}{2}|V|$, we have
    $|E(\Omega,\Omega^{c})|\geq|\Omega|.$
    If we further assume $(\alpha,\beta)\neq(2,2)$, then
    $|E(\Omega,\Omega^{c})|\geq\left(1+\max\left\{\frac{\alpha}{2},\left\lceil\frac{(\beta-\alpha)\alpha}{\beta-1}\right\rceil\right\}\right)|\Omega|.$
  • (ii)
    Suppose $2d(\beta-2)\geq(\alpha-\beta)(\alpha-4\beta)$. Then for any subset $\Omega\subseteq V$, we have
    $\min_{L\sqcup R=\Omega}\left(|E(L,L)|+|E(R,R)|\right)+|E(\Omega,\Omega^{c})|\geq|\Omega|.$
    where the minimization is taking over all possible partitions $\{L,R\}$ of $\Omega$, where one of $L$ and $R$ can be empty.
Proof.
This follows directly from the fact $d-\theta_{n-1}\leq 2dh$ and $d+\theta_{1}\leq 2d\beta_{0}$ and the eigenvalue estimates proved in Subsection 5.1. ∎
Now, we recall the definition of expander.
Definition 5.8 ([2, p.151]).
A graph $G=(V,E)$ is called an ($n,d,c$)-expander if it has $n$ vertices, the maximum degree of a vertex is $d$, and, for every set of vertices $W\subset V$of cardinality $|W|\leq n/2$, the inequality $|N(W)|\geq c|W|$ holds, where $N(W)$ denotes the set of all vertices in $V\backslash W$ adjacent to some vertex in $W$.
The following theorem mentioned in [2, p.153] shows a correspondence between the second largest eigenvalue of adjacency matrix of a graph and its expansion properties. For a proof, we refer to [1, Theorem 4.3].
Theorem 5.9 ([2, p.153]).
A $d$-regular graph $G$ with $n$ vertices is an ($n$, $d$, $c$)-expander for $c=2(d-\theta_{n-1})/(3d-2\theta_{n-1})$.
Combining with our eigenvalue bound, we obtain the following conclusion.
Theorem 5.10.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. If $1\neq\beta\geq\alpha$ and $(\alpha,\beta)\neq(2,2)$, then $G$ is an ($n$, $d$, $c$)-expander for
$c=1-\frac{d}{d+4+\max\left\{\alpha,2\left\lceil\frac{(\beta-\alpha)\alpha}{\beta-1}\right\rceil\right\}}.$
That is, for any $W\subset V$ with $|W|\leq n/2$, we have
$|N(W)|\geq\left(1-\frac{d}{d+4+\max\left\{\alpha,2\left\lceil\frac{(\beta-\alpha)\alpha}{\beta-1}\right\rceil\right\}}\right)|W|.$
Proof.
It is a direct consequence of Theorem 1.6 and Theorem 5.9 . ∎
5.3 Application in bounding the volume growth
Theorem 5.11.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $\kappa_{LLY}(x,y)\geq k>0$ for any $xy\in E$. Then, we have for any $x\in V$ and $i\geq 1$ that
$|S_{i+1}(x)|\leq\frac{d-\max\left\{\alpha+1,ikd/2\right\}}{\beta}|S_{i}(x)|.$
We need the following Lemma shown by Lin, Lu and Yau [25, Lemma 4.4]. Let $x$ be a given vertex. For any $y\in S_{i}(x)$, $i\geq 1$, we use the following notations:
$d^{x,+}_{y}:=|\{z\in S_{i+1}(x)|\,yz\in E\}|\,\,\text{and}\,\,d^{x,-}_{y}:=|\{z\in S_{i-1}(x)|\,zy\in E\}|.$
Lemma 5.12.
For any two distinct vertices $x$ and $y$, we have
$\kappa_{LLY}(x,y)\leq\frac{1+(d^{x,-}_{y}-d^{x,+}_{y})/d_{y}}{d(x,y)}.$
Proof of Theorem 5.11.
Let $x$ be a given vertex. For any $y\in S_{i}(x)$ with $i\geq 1$, there exists a vertex $u\in S_{i-1}(x)$ such that $uy\in E$, and hence $u$ and $y$ share $\alpha$ common neighbors. These common neighbors can not fall into $S_{i+1}(x)$. Hence, we deduce that
$d^{x,+}_{y}\leq d-\alpha-1.$ (5.1)
On the other hand, for any $w\in S_{i+1}(x)$ with $i\geq 1$, there exists a vertex $v\in S_{i-1}(x)$ such that $d(v,w)=2$. The two vertices $v$ and $w$ share $\beta$ common neighbors. These common neighbors all lie in $S_{i}(x)$. Hence we have
$d^{x,-}_{w}\geq\beta.$ (5.2)
Therefore, we estimate from $\sum_{w\in S_{i+1}(x)}d_{w}^{x,-}=E(S_{i}(x),S_{i+1}(x))=\sum_{y\in S_{i}(x)}d_{y}^{x,+}$ that
$|S_{i+1}(x)|\leq\frac{d-\alpha-1}{\beta}|S_{i}(x)|.$ (5.3)
Due to our curvature condition, it follows by Lemma 5.12 that $idk\leq d+d^{x,-}_{y}-d^{x,+}_{y}$. Since $d^{x,-}_{y}\leq d-d^{x,+}_{y}$, we have
$d^{x,+}_{y}\leq\left(1-\frac{ik}{2}\right)d.$ (5.4)
Hence, by a similar argument as in (5.3), we derive
$|S_{i+1}(x)|\leq\frac{\left(1-\frac{ik}{2}\right)d}{\beta}|S_{i}(x)|.$ (5.5)
Inequalities (5.3) and (5.5) together yield the desired result. ∎
Remark 5.13.
From the proof, it is straightforward to show that Theorem 5.11 holds still for graphs with maximal vertex degree $d$ such that any two adjacent vertices have at least $\alpha$ common neighbors and any two vertices at distance $2$ have at least $\beta$ common neighbors. A volume estimate for this class of graphs with diameter $2$ has been obtained by Seidel [40, Page 158] (see also [10, Proposition 1.4.1]). Our results can be considered as a generalization of Seidel’s volume estimate.
To conclude this subsection, we prove Theorem 1.8 from the Introduction. We need the following key lemma due to Terwilliger [46, Lemma 3.3] about $(s,c,a,k)$-graphs. Below we state a particular case of [46, Lemma 3.3] that $s=2$, i.e., the case of $(2,c,a,k)$-graphs, which are amply regular, see also [10, Proposition 1.9.1].
Lemma 5.14.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $1\neq\beta\geq\alpha$. For any $x\in V$ and $w\in S_{i+1}(x)$ with $i\geq 1$, we have
$d_{w}^{x,-}\geq\beta+i-1.$
Proof of Theorem 1.8.
The proof is done by applying the procedure in the proof for Theorem 5.11 via replacing (5.2) by
$d_{w}^{x,-}\geq\beta+i-1,\,\,\text{for}\,\,w\in S_{i+1}(x)$ (5.6)
and replacing (5.4) by
$d_{y}^{x,+}\leq d-i\left(1+\frac{1}{2}\left\lceil\frac{\alpha(\beta-\alpha)}{\beta-1}\right\rceil\right),\,\,\text{for}\,\,y\in S_{i}(x).$ (5.7)
Notice that in (5.6) and (5.7) we have applied Lemma 5.14 and Theorem 3.1, respectively. ∎
In fact, Theorem 1.8 provides sharp volume estimate. The equality can be achieved for hypercube graphs.
Example 5.15.
For any $d\geq 2$, the $d$-dimensional hypercube graph $Q^{d}$ is amply regular with parameters $(2^{d},d,0,2)$. Hence it satisfies the restriction $1\neq\beta\geq\alpha$ in Theorem 1.8. Moreover, the Lin–Lu–Yau curvature of each edge in $Q^{d}$ is equal to $\frac{2}{d}$. Applying Theorem 1.8 to $Q^{d}$ yields the estimates
$|S_{i+1}(x)|\leq\frac{d-i}{i+1}|S_{i}(x)|,\,\,\text{for}\,\,i\geq 1.$ (5.8)
On the other hand, we have $|S_{i}(x)|=\binom{d}{i}$ for $1\leq i\leq\mathrm{diam}(Q^{d})=d$. That is, all the inequalities in (5.8) are equalities. Therefore, the estimate for $|G|=\sum_{i\geq 0}|S_{i}(x)|$ derived from Theorem 1.8 is also sharp for $Q^{d}$.
Acknowledgement
This work is supported by the National Key R & D Program of China 2020YFA0713100 and the National Natural Science Foundation of China No. 12031017 and No. 12431004. We are very grateful to Shuliang Bai, Jack H. Koolen, Qianqian Yang for discussions on explicit constructions of amply regular graphs, especially in the case $\beta=1$.
Appendix
In this appendix, we construct a sequence of amply regular graphs $\{G_{n}\}$ with parameters $\left(3(2^{n+1}-1),4,1,1\right)$ such that ${\rm diam}(G_{n})$ tends to infinity.
We first construct a sequence of graphs $\{H_{n}\}$. Let $H_{0}$ be a triangle. For $k\geq 0$, we construct $H_{k+1}$ based on $H_{k}$ as follows: Let $\{v_{1},...,v_{s}\}$ be the set of vertices of degree $2$ in $H_{k}$. Add $2s$ number of vertices $\{v_{1a},v_{1b},...,v_{sa},v_{sb}\}$. For $1\leq i\leq s$, connect $v_{i}v_{ia}$, $v_{i}v_{ib}$ and $v_{ia}v_{ib}$. Then, we obtain the graph $H_{k+1}$.
Figure 1: The construction of $H_{0}$, $H_{1}$ and $H_{2}$
Now, we construct $G_{n}$ based on $H_{n}$ for $n\geq 2$. We can check by induction that there are $3\times 2^{n}$ number of vertices of degree $2$ in $H_{n}$. We label them every $12$ vertices a group as follows $1,2,3,4,1,3,2,4,1,4,2,3,5,6,7,8,5,7,6,8,5,8,6,7,...$(see Figure 2). Connecting every two vertices with the same label, we obtain the graph $G_{n}$.
Figure 2: The construction of $G_{3}$
We can directly check that, for $n\geq 2$, $G_{n}$ is an amply regular graph with parameters $\left(3(2^{n+1}-1),4,1,1\right)$. Considering the distance between the first labeled vertex and the last labeled vertex, we deduce that the diameter of $G_{n}$ is $12n+$ 1.